A Metal oxide contains 20% oxygen by weight .  Find the Equivalent weight of metal .

Grade:11

2 Answers

Khimraj
5 years ago
metal = 80% 
oxygen = 20%

100g of metaloxide contains 80g metal and 20g oxygen,therefore

equivalent weight of metal is=mass of metal x 8/mass of oxygen = (80/20)x 8=32

Hope it clears.

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
Let us consider 100 g of Metal oxide.
Mass of metal in oxide = 80 g
Mass of oxygen in oxide = 20 g
Now, equivalents of Metal = equivalents of oxygen
Hence, 80/x = 20/8
Hence, x = 4 x 8 = 32 g
Hence, Equivalent weight of Metal = 32 g
 
Thanks and regards,
Kushagra